# Pulling apart the mechanisms that lead to jammed knitted fabrics

### v1

@@ VII The Jamming Mechanism
Using the simulations, we are able to identify four dominant types of contacts for a single stitch: (1) self-contact where yarn within a stitch is in contact with a different segment of yarn within the same stitch, (2) nearest-neighbor contact with stitches in the same row but different columns, (3) nearest-neighbor contact with stitches in the same column a row above or below, and (4) next-nearest-neighbor contact with stitches in the same column two rows above or below. Each of these contacts are shown individually in Fig. 6a-d and together in Fig. 6e,f. We consider the forces applied by each of the yarn-yarn contacts (discussed more in Supplementary Section V and Figs. S9, S10, and S11).
@@
When stretched in the $x$-direction, jammed fabrics must first rearrange out-of-plane before they can extend. During conventional linear elastic behavior of unjammed fabrics, the stitch enlarges in the $z$-direction. By contrast, jammed fabrics shrink in the $z$-direction until they reach the softened regime and then expand, as shown in Fig. 6g and Fig S12. The only contact applying non-negligible force is the nearest-neighbor same-column contact, but this contact is applying a force that would expand the stitch out-of-plane like unjammed fabrics, not contract as is seen in jamming (Fig. S11). Since none of the contacts are applying an out-of-plane contraction force, this indicates that the yarn bending is causing this motion, in agreement with the energy gap analysis.
@@
For jamming in fabrics pulled in the $y$-direction, the energy gap data indicates that yarn contacts lead to jammed behavior. Analysing the individual contact areas shows that it is the the interplay between contacts that leads to changing mechanics. For un-jammed fabrics, the contact forces in the $y$-direction dominate in the stress-free configuration; the second-column contact force is the largest contributing contact force and it pushes the stitch to extend in the $y$-direction (see Fig. S11e). For jammed fabrics, the contact forces in the $x$-direction are larger than those in the extension direction. The sum of the self contacts and same row contacts, both of which promote the stitch to expand in the $x$-direction (Fig. S11d), become larger than the second-column and same-column contacts in the $y$-direction, shown in Fig. 6h. These forces inhibit the fabric from the classical, in-plane Poisson response and prevent extension in the $y$-direction from occurring without a high enough applied force to overcome these initial, internal forces.
@@
same-column contact - compress
second same-column contact - extend
Table 2: Summary of the direction of the contact force at each contact area at zero stress.
@@ VIII Conclusion
Knitted fabric jamming is an initial region of low-stress high modulus behavior that can be seen in experimental samples. As we decrease the “stitch size” in knitted fabric manufacturing, fabrics become more jammed as described by the secant modulus $Y_{\rm sec}$ and the offset stress $\sigma_{0}$, two metrics we have established to quantify jamming. Using these metrics, we compare experimental results to simulated changes to the length of yarn per stitch and yarn radius. Error in experimental yarn radius measurements makes quantification difficult, but the simulation results are qualitatively consistent with experiment. These geometrically-confined jammed states are energy minima for the fabrics, but are generally not contact energy minima (for fabrics jammed in the $x$-direction) or bending energy minima (for fabrics jammed in the $y$-direction). The energy difference between the lowest energy state and the lowest energy for either bending or contact energy shows that there is a soft transition region that is dependent on the compressibility of the yarn. Analyzing the shape of the fabric shows anisotropic markers where jamming in the $x$-direction can be seen in the ratio of the fabric length to fabric width and jamming in the $y$-direction can be found in the area strain. Though we cannot accurately measure the out-of-plane thickness of experimental samples, simulations suggest that fabrics jammed in the $x$-direction first contract in the $z$-direction before expanding, and that this behavior is not caused by yarn contacts but instead by yarn rearrangements to reduce yarn bending. Contact forces show that all four primary contacts contribute to knitted fabric jamming in the $y$-direction by inhibiting linear elastic contraction in the $x$-direction.
@@ Acknowledgements.
Each fabric was manufactured on a STOLL industrial knitting machine using the proprietary M1 software and are 64 stitches wide and 64 stitches tall without buffers. The yarn is Tamm Petit 2/30 T4201 White 100% acrylic yarn from The Knit Knack Shop™. The samples are stockinette fabric, also known as jersey or plain knit. Samples were made at different machine tensions settings, 10.5-13. Smaller and larger tension settings were not able to produce consistent samples without holes or machine error. For each individual stretching direction, a sample was made with an additional buffer region in the direction of extension. The buffer region was clamped during extension to ensure that the entire sample is free to extend during the experimental procedure.
@@
S2 Machine Tension Settings
The knitting machine does not have a direct input for the length of yarn per stitch of the fabric. Instead, the manufacturer programs in a machine tension. This tension affects the yarn per stitch and leads to a tighter (smaller tension value) or looser (larger tension value) fabric, shown in Fig. S2a. The length of yarn per stitch is measured by making fabric samples without the buffer region, with side lengths of 32, 48, and 64 stitches. The samples are weighed on a scale with 0.01 g precision, and we then use the length to weight ratio of the yarn provided by the yarn manufacturer to calculate an average length of yarn per stitch. The machine tension setting also affects the radius of the yarn within the fabric, shown in Fig. S2b. We measure the yarn radius in situ by taking a picture of each fabric sample at rest on the tabletop and measuring the average yarn radius with a digital length and pixel conversion for fifteen different locations within each fabric. As a consequence of changing length of yarn per stitch and yarn radius, the size of the stitch cell also changes, shown in Fig. S2c,d. The stitch cell dimensions where taken using a caliper with 0.01 mm precision on samples made without the buffer regions. Each sample was measured five times and these results are given in S2.
@@
Figure S2: How the (a) length of yarn per stitch, (b) yarn radius, (c) stitch cell dimensions, and (d) stitch cell area changes with the machine tension setting on samples without buffers.
@@
As discussed in Section S2, changing the machine tension setting on the industrial knitting machine is confirmed to change the length of yarn per stitch $L$ and the yarn radius $r$. What happens to the core radius $r_{core}$ is yet unknown and will require further study. As these parameters change, the jamming metrics $Y_{\rm sec}$ and $\sigma_{0}$ also change, as shown in Fig. 2e,f in the Main Text for the x-direction. Fig. S4 shows how these metrics change when the fabric is stretched in the y-direction. Fits and fitting parameters are given in 3. All fits were completed using the NonlinearModelFit[34] function in Mathematica version 14.0. Multiple forms were attempted for each curve and the form with the largest $R^{2}$ value was selected.
Figure S4: Secant modulus $Y_{\rm sec}$ (a) and offset stress $\sigma_{0}$ (b) for changing length of yarn per stitch $L$ (green), yarn radius $r$ (magenta), and yarn radius $r$ for fixed ratio $r_{core}/r=0.45$ (blue). The length of yarn per stitch is $L=12.2$ mm for the yarn radius sweep and $L=12.0$ mm for the yarn radius with fixed $r_{core}/r$ sweep. For the fixed ratio $r_{core}/r$ radius sweep, all values of the yarn radius attempted yielded fabric that is jammed in the y-direction.
@@
y Experiment $Y_{\rm sec}$ $(1*10^{9}\pm 7*10^{9})e^{(-1.1\pm 0.3)L/r}$ 0.9472
y Experiment $\sigma_{0}$ $(6*10^{6}\pm 1*10^{8})e^{(-1\pm 1)L/r}$ 0.6941
@@
Figure S5: Secant modulus, $Y_{\rm sec}$, of the stress-strain curves for fabrics of different (a) compression scaling constant $k$, (c) bending modulus $B$, and (e) core radii $r_{core}$ pulled in both the x- (magenta) and y-directions (cyan). Offset stress, $\sigma_{0}$, of the stress-strain curves for fabrics of different (b) compression scaling constant $k$, (d) bending modulus $B$, and (f) core radii $r_{core}$ pulled in both the x- (magenta) and y-directions (cyan).
Figure S6: Onset of unjamming $L_{trans}$ as a function of bending modulus for four different values of the compression scaling constant $k$: (a) 0.1 $mN/mm^{2}$, (b) 0.5 $mN/mm^{2}$, (c) 1.0 $mN/mm^{2}$, and (d) 1.7 $mN/mm^{2}$. These plots were used to make the composite figure in Main Text Fig. 4c, where $r=0.74mm$.
@@
Fabrics can become so stiff that the linear elastic regime seems to disappear entirely and the fabric moves smoothly from jamming to non-linear strain-stiffening. This state can be reached by varying any of the simulation parameters sufficiently. For example, Fig. S7 shows how increasing the compression scaling constant $k$ can reduce and then eliminate the linear elastic regime for stiff fabrics. Fig. S8 shows how the jamming metrics change with the increasing compression scaling constant. As seen with the transition from jammed to unjammed, there is a measurable difference in both jamming metrics for these different mechanical regimes. Moving directly from unjammed fabrics to this very stiff regime in phase space can thus obscure the jamming transition in the jamming metric data. This is particularly true for the offset stress, $\sigma_{0}$. If analyzing data algorithmicly with the same bounds for calculating $\sigma_{0}$, you can observe decreasing $\sigma_{0}$ as you move into this jammed-to-nonlinear-strain-stiffening regime, as seen in the y-extension data for the compression scaling constant sweep data shown in Fig. S5b. This is particularly relevant for extension in the y-direction, which is generally stiffer [15].
Figure S7: Stress-strain plots for simulations run for different values of the compression scaling constant, $k$, detailed on the right. These fabrics were stretched in the x-direction and had a length of yarn per stitch of $L=10.7$ mm.
Figure S8: Jamming metric data for the stress-strain curves shown in Fig. S7, all of which are jammed and some of which have little to no perceivable linear elastic regime between jamming and non-linear strain stiffening. (a) Secant modulus, $Y_{\rm sec}$, of the stress-strain curves for fabrics of different compression scaling constants pulled in the x-direction. (b) Offset stress, $\sigma_{0}$, of the stress-strain curves for fabrics of different compression scaling constants pulled in the x-direction. The length of yarn per stitch is $L=10.7$ mm. Other simulation parameters are as described in Main Text Table 1.
@@
We calculated the packing fraction as a function of strain for the set of yarn lengths $L\in[10.5,...,12.2]$ mm for $k=0.1$ $mN/mm^{2}$, shown in Fig. S9. From this data, there is no characteristic packing fraction that indicates jamming. Sweeping over yarn length for zero strain, the packing fraction that indicates the end of jamming is $\phi\approx 0.85$ for this set of yarn parameters, listed in Main Text Table 1. However, sweeping out in strain for sample $L=10.5$mm, $\phi=0.85$ is at $(\epsilon_{x},\epsilon_{y})$ location $(0.325,0.177)$, which is far beyond the jamming transition in both extension directions and is even contained within the non-linear strain-stiffening regime. This indicates that any critcal packing fraction in the $L$ phase space is not necessarily critical in strain. Discontinuities in the packing fraction data make additional analysis, such as transition packing fraction along the strain sweeps as a function of yarn length, unreliable. These jumps are not the result of the voxel size, shown in S10b. Though there is a maximum voxel side length for reliable volume data (see Fig. S10a), significantly reducing this voxel side length did not result in smoother data. We believe these jumps are the result of small local rearrangements within the stitch. Since the contact potential is very non-linear and highy dependent on the depth of the contact, the compressed volume of the stitch does not meaningfully represent the energy being minimized in the simulation. The total volume of the stitch can vary widely without subsequent jumps in the energy of the stitch. It is possible that a dynamic simulation framework would result in smoother packing fraction data.
Figure S9: Packing fraction as a function of strain for different lengths of yarn per stitch, $L$. Fabrics were extended in both the x- (left) and y-directions (right). Fabrics that display jammed behavior have a length of yarn per stitch $L\leq 11.5$ mm. For these simulations, the compression scaling constant is $k=0.1mN/mm^{2}$.
Figure S10: (a) Calculated total volume of the stitch within the fabric using the voxel method as a function of voxel side length. Small voxel side lengths ($<0.05$) show stability in the calculated compressed volume of the stitch, whereas larger voxel side lengths show wide variation in volume. For these simulations, the length of yarn per stitch is $L=12.2mm$ and the compression scaling constant is $k=0.1mN/mm^{2}$. (b) Packing fraction as a function of strain for fabric stretched in the x-direction for two different voxel side lengths: $l=0.04$ in blue and $l=0.02$ in black. For these simulations, the length of yarn per stitch is $L=10.5mm$ and the compression scaling constant is $k=0.1mN/mm^{2}$.
@@
Each contact volume exerts a force on the yarn within the stitch. Individual forces for each contact for a single set of yarn and manufacturing parameters are shown in Fig. S11. The only non-negligible contact forces in the x-direction are the nearest-neighbor same row contacts and the self-contact. Both of these forces are reduced when the fabric extends in the x-direction (Fig. S11a), and thus assist with normal linear elastic behavior when the fabric is pulled in the x-direction and inhibit it when the fabric is pulled in the y-direction (Fig. S11d). The nearest-neighbor same column contact and the second-nearest-neighbor same column contact are applying forces in the y-direction; the nearest-neighbor same column contact encourages the fabric to shrink in the y-direction while the second-nearest-neighbor same column contact encourages the fabric to grow in the y-direction. These forces are stagnant at low strains when the fabric is pulled in the x-direction, potentially indicating a lack of contact rearrangement. When the fabric is pulled in the y-direction, the force from the second-nearest-neighbor same column contact decreases in magnitude while the nearest-neighbor same column contact grows. The interplay between these two forces and the forces in the x-direction determines jamming in the y-direction, as discussed in the Main Text and displayed in Fig. 6h.
@@
When the fabric is stretched in the x-direction, the mechanics of jamming show the stitch shrinking out-of-plane (Main Text Fig. 6g). The contact with non-negligible force in this direction is the nearest-neighbor same column contact, which is applying a force in the direction of expansion, directly contrary to the actual motion of the stitch. This is additional evidence that bending forces dominate jamming in the x-direction.
@@
Figure S11: Contact forces as a function of strain in each of the relevant directions for all of the contacts. The sign of the force indicates whether or not it assists with normal linear elastic behavior (positive) or resists linear elastic mechanics (negative) when stretched in that direction. For each type of contact where relevant, both contacts of that type are added (eg: the same column contact is the sum of this contact force with the stitch above and below). Both sets of data are for length of yarn per stitch $L=10.5$ mm, which is jammed in both directions. (a) Force in the x-direction for the fabric stretched in the x-direction. The only non-negligible forces are from the same row and self contacts, which track together. (b) Force in the y-direction for the fabric stretched in the x-direction. The nearest-neighbor same column contact promotes typical linear elastic extension while the second-nearest neighbor contact resists that motion. (c) Force in the z-direction (out-of-plane) for the fabric stretched in the x-direction. The only contribution in this direction is the same column contact, which is pushing out, contrary to the actual motion of the yarn in this jammed configuration. (d) Force in the x-direction for fabric stretched in the y-direction. The same row and self contacts are the only contributing contacts, and both are in the direction of expansion. (e) Force in the y-direction for fabric stretched in the y-direction. The only contributions are the nearest-neighbor same column and second-nearest-neighbor same column contacts, which resist linear elastic motion and promote it, respectively. The forces in the z-direction for fabrics stretched in the y-direction are similar to those found when stretching in the x-direction, but flipped in sign due to the different out-of-plane linear elastic behavior seen in the y-direction (Fig. S12b).
@@
As discussed in the main text, jammed fabrics pulled in the x-direction have unique mechanics markers in the out-of-plane (z) strain. Unjammed fabrics have negative Poisson ratio at low strain and jammed fabrics have positive Poisson ratio at low strain. Fig. S12 shows the full range of motion for one unjammed sample and a superjammed sample. In the superjammed sample, the Poisson ratio never changes sign through the entirety of the range of motion. This is due to a lack of a linear elastic regime, which is discussed in more detail in Section S4.
@@
For simulated fabrics pulled in the x-direction, the stitch cell area strain is non-monotonic, first increasing and then beginning to decrease at very large strain as seen in S13. Unlike fabrics stretched in the y-direction, there is no marker of jamming in the simulated area strain for fabrics stretched in the y-direction. Though we cannot directly compare these results to experimental area strain, seen in S14, because of the machine tension setting affect on yarn radius and potentially the core radius, both the simulations and experiments show periods of increasing area strain followed by decreasing area strain. The simulations also do not consider boundary effects, which would provide large constraints on the cell area, particularly at high strain. These boundary effects may be the cause of the relatively constant area strain at large uniaxial strain, which is not present in simulations.
@@
extension direction $k$ [mNmm-2] $L$ [mm] Step Size [mm] $\epsilon_{final}$ RMS [J]

### v2

@@ VII The Jamming Mechanism
Using the simulations, we are able to identify four dominant types of contacts for a single stitch: (1) self-contact where yarn within a stitch is in contact with a different segment of yarn within the same stitch, (2) nearest-neighbor contact with stitches in the same row but different columns, (3) nearest-neighbor (NN) contact with stitches in the same column a row above or below, and (4) next-nearest-neighbor (NNN) contact with stitches in the same column two rows above or below. Each of these contacts are shown individually in Fig. 6a-d and together in Fig. 6e,f. We analyze the forces applied by each of the yarn-yarn contacts in Fig. 6f and Supplementary Section V (Figs. S9, S10, and S11).
@@
When stretched in the $x$-direction, jammed fabrics must first rearrange out-of-plane before they can extend. During conventional linear elastic behavior of unjammed fabrics, the stitch enlarges in the $z$-direction. By contrast, jammed fabrics shrink in the $z$-direction until they reach the softened regime and then expand, as shown in Fig. 6g and Fig S12. The only contact applying non-negligible force is the NN same-column contact, but this contact is applying a force that would expand the stitch out-of-plane like unjammed fabrics, not contract as is seen in jamming (Fig. S11). Since none of the contacts are applying an out-of-plane contraction force, this indicates that the yarn bending is causing this motion, in agreement with the energy gap analysis.
Figure 6: (a-e) Stitch contact volumes shown for a jammed fabric with a length of yarn per stitch $L=10.5~{}{\rm mm}$. (a) The self-contact shown in green. (b) The nearest-neighbor contact with stitches in the same row but different columns shown in magenta. (c) The nearest-neighbor contact with stitches in the same column a row above (lavender) or below (cyan). (d) The next-nearest-neighbor contact with stitches in the same column two rows above or below in yellow. (e) All contact volumes shown together in the stitch for a jammed fabric with length of yarn per stitch $L=10.5~{}{\rm mm}$. (f) Contact volumes shown for an un-jammed fabric with length of yarn per stitch $L=12.2~{}{\rm mm}$. The contact volumes for (a-f) look fuzzy at the edges because they are made of a semi-transparent clouds of points. (g) Strain in the $z$-direction as a function of strain in the $x$-direction for an un-jammed ($L=12.2~{}{\rm mm}$ in filled circles) and jammed ($L=10.5~{}{\rm mm}$ in open circles) fabric being stretched in the $x$-direction. The grey horizontal line is zero-strain. (h) Magnitude of the ratio of total contact forces in the $x$- and $y$-directions for zero-strain configurations of fabrics with different lengths of yarn per stitch $L$. Near the jamming transition, we see that the contact force in the $y$-direction becomes larger than the contact force in the $x$-direction. The dashed horizontal line represents equal contact forces in the $x$- and $y$-directions. Error in contact force is described in Table S3.
@@
For jamming in fabrics pulled in the $y$-direction, the energy gap data indicates that yarn contacts lead to jammed behavior. Analyzing the individual contact areas shows that it is the the interplay between contacts that leads to changing mechanics. For un-jammed fabrics, the contact forces in the $y$-direction dominate in the stress-free configuration; the NNN same-column contact force is the largest contributing contact force and it pushes the stitch to extend in the $y$-direction (see Fig. S11e). For jammed fabrics, the contact forces in the $x$-direction are larger than those in the extension direction. The sum of the self contacts and same-row contacts, both of which expand the stitch in the $x$-direction (Fig. S11d), become larger than the NNN same-column and NN same-column contacts in the $y$-direction, shown in Fig. 6h. These forces inhibit the fabric from the classical, in-plane Poisson response and prevent extension in the $y$-direction from occurring without a high enough applied force to overcome these initial, internal forces.
@@
NN same-column contact - compress
NNN same-column contact - extend
Table 2: Summary of the direction of the contact force at each contact area at zero strain.
@@ VIII Conclusion
Knitted fabric jamming corresponds to an anomalously-stiff regime of fabric response at low strain that can be seen in experimental samples. As we decrease the “stitch size” in knitted fabric manufacturing, fabrics become more jammed as described by the secant modulus $Y_{\rm sec}$ and the offset stress $\sigma_{0}$, two metrics we have established to quantify jamming. Using these metrics, we compare experimental results to simulated changes to the length of yarn per stitch and yarn radius. The simulation results are qualitatively consistent with experiment and agree that fabrics are “more jammed” in the $y$-direction. These geometrically-confined jammed states are energy minima for the fabrics, but are generally not contact energy minima (for fabrics jammed in the $x$-direction) or bending energy minima (for fabrics jammed in the $y$-direction). This indicates that these states are characterized by distinct distributions of residual stress. The energy difference between the zero-strain state and the lowest energy for either bending or contact energy shows that there is a soft transition region that is dependent on the compressibility of the yarn. Analyzing the shape of the fabric shows anisotropic markers where jamming in the $x$-direction can be seen in the aspect ratio of the stitches and jamming in the $y$-direction can be found in the area strain. Though we cannot accurately measure the out-of-plane thickness of experimental samples, simulations suggest that fabrics jammed in the $x$-direction first contract in the $z$-direction before expanding, and that this behavior is not caused by yarn contacts but instead by yarn rearrangements to reduce yarn bending. Contact forces show that all four primary contacts contribute to knitted fabric jamming in the $y$-direction by inhibiting linear elastic contraction in the $x$-direction.
@@ Acknowledgements.
Each fabric was manufactured on a STOLL industrial knitting machine using the proprietary M1 Plus software and are 64 stitches wide and 64 stitches tall without buffers. The yarn is Tamm Petit 2/30 T4201 White 100% acrylic yarn from The Knit Knack Shop™. The samples are stockinette fabric, also known as jersey or plain knit. Samples were made at different “stitch size” settings, 10.5-13. Smaller and larger “stitch size” settings were not able to produce consistent samples without holes or machine error. For each individual stretching direction, a sample was made with an additional buffer region in the direction of extension. The buffer region was clamped during extension to ensure that the entire sample is free to extend during the experimental procedure.
@@
S2 “Stitch Size”
The knitting machine does not have a direct input for the length of yarn per stitch of the fabric. Instead, the manufacturer indirectly programs in a machine tension via the “stitch size” parameter. This tension affects the yarn per stitch and leads to a tighter (smaller “stitch size”) or looser (larger “stitch size”) fabric, shown in Fig. S2a. The length of yarn per stitch is measured by making fabric samples without the buffer region, with side lengths of 32, 48, and 64 stitches. The samples are weighed on a scale with 0.01 g precision, and we then use the length to weight ratio of the yarn provided by the yarn manufacturer to calculate an average length of yarn per stitch. The “stitch size” also affects the radius of the yarn within the fabric, shown in Fig. S2b. We measure the yarn radius in situ by taking a picture of each fabric sample at rest on the tabletop and measuring the average yarn radius with a digital length and pixel conversion for fifteen different locations within each fabric. As a consequence of changing length of yarn per stitch and yarn radius, the size of the stitch cell also changes, shown in Fig. S2c,d. The stitch cell dimensions where taken using a caliper with 0.01 mm precision on samples made without the buffer regions. Each sample was measured five times and these results are given in S2.
@@
Figure S2: How the (a) length of yarn per stitch, (b) yarn radius, (c) stitch cell dimensions, and (d) area-per-stitch change with the “stitch size” setting on samples without buffers. Error bars represent the mean standard error.
@@
As discussed in Section S2, changing the machine tension setting via the “stitch size” on the industrial knitting machine is confirmed to change the length of yarn per stitch $L$ and the yarn radius $r$. What happens to the core radius $r_{core}$ is yet unknown and will require further study. As these parameters change, the jamming metrics $Y_{\rm sec}$ and $\sigma_{0}$ also change, as shown in Fig. 2e,f in the Main Text for the $x$-direction. Fig. S4 shows how these metrics change when the fabric is stretched in the $y$-direction. Fits and fitting parameters are given in 3. All fits were completed using the NonlinearModelFit[34] function in Mathematica version 14.0. Multiple forms were attempted for each curve and the form with the largest $R^{2}$ value was selected.
Figure S4: Secant modulus $Y_{\rm sec}$ (a) and offset stress $\sigma_{0}$ (b) for changing length of yarn per stitch $L$ (gray), yarn radius $r$ (orange), and yarn radius $r$ for fixed ratio $r_{core}/r=0.45$ (blue) for fabrics stretched in the $y$-direction. The length of yarn per stitch is $L=12.2~{}{\rm mm}$ for the yarn radius sweep and $L=12.0~{}{\rm mm}$ for the yarn radius with fixed $r_{core}/r$ sweep. For the fixed ratio $r_{core}/r$ radius sweep, all values of the yarn radius attempted yielded fabric that is jammed in the $y$-direction.
@@
y Experiment $Y_{\rm sec}$ $(4*10^{8}\pm 2*10^{9})e^{(-1.0\pm 0.2)L/r}$ 0.9716
y Experiment $\sigma_{0}$ $(4*10^{3}\pm 2*10^{4})e^{(-0.7\pm 0.2)L/r}$ 0.9508
@@
Figure S5: Secant modulus, $Y_{\rm sec}$, of the stress-strain curves for fabrics of different (a) compression scaling constant $k$, (c) bending modulus $B$, and (e) core radii $r_{core}$ pulled in both the $x$- (purple) and $y$-directions (blue). Offset stress, $\sigma_{0}$, of the stress-strain curves for fabrics of different (b) compression scaling constant $k$, (d) bending modulus $B$, and (f) core radii $r_{core}$ pulled in both the $x$- (purple) and $y$-directions (blue).
Figure S6: Onset of unjamming $L_{trans}$ as a function of bending modulus for four different values of the compression scaling constant $k$: (a) 0.1 ${\rm mN/mm^{2}}$, (b) 0.5 ${\rm mN/mm^{2}}$, (c) 1.0 ${\rm mN/mm^{2}}$, and (d) 1.7 ${\rm mN/mm^{2}}$. These plots were used to make the composite figure in Main Text Fig. 4c, where $r=0.74~{}{\rm mm}$.
@@
Fabrics can become so stiff that the linear elastic regime seems to disappear entirely and the fabric moves smoothly from jamming to non-linear strain-stiffening. This state can be reached by varying any of the simulation parameters sufficiently. For example, Fig. S7 shows how increasing the compression scaling constant $k$ can reduce and then eliminate the linear elastic regime for stiff fabrics. Fig. S8 shows how the jamming metrics change with the increasing compression scaling constant. As seen with the transition from jammed to unjammed, there is a measurable difference in both jamming metrics for these different mechanical regimes. Moving directly from unjammed fabrics to this very stiff regime in phase space can thus obscure the jamming transition in the jamming metric data. This is particularly true for the offset stress, $\sigma_{0}$. If analyzing data algorithmically with the same bounds for calculating $\sigma_{0}$, you can observe decreasing $\sigma_{0}$ as you move into this jammed-to-nonlinear-strain-stiffening regime, as seen in the $y$-extension data for the compression scaling constant sweep data shown in Fig. S5b. This is particularly relevant for extension in the $y$-direction, which is generally stiffer [15].
Figure S7: Stress-strain plots for simulations run for different values of the compression scaling constant, $k$, detailed on the right. These fabrics were stretched in the $x$-direction and had a length of yarn per stitch of $L=10.7~{}{\rm mm}$.
Figure S8: Jamming metric data for the stress-strain curves shown in Fig. S7, all of which are jammed and some of which have little to no perceivable linear elastic regime between jamming and non-linear strain stiffening. (a) Secant modulus, $Y_{\rm sec}$, of the stress-strain curves for fabrics of different compression scaling constants pulled in the $x$-direction. (b) Offset stress, $\sigma_{0}$, of the stress-strain curves for fabrics of different compression scaling constants pulled in the $x$-direction. The length of yarn per stitch is $L=10.7~{}{\rm mm}$. Other simulation parameters are as described in Main Text Table 1.
@@
We calculated the packing fraction as a function of strain for the set of yarn lengths $L\in[10.5,...,12.2]$ mm for $k=0.1$ $mN/mm^{2}$, shown in Fig. S9. From this data, there is no characteristic packing fraction that indicates jamming. Sweeping over yarn length for zero strain, the packing fraction that indicates the end of jamming is $\phi\approx 0.85$ for this set of yarn parameters, listed in Main Text Table 1. However, sweeping out in strain for sample $L=10.5~{}{\rm mm}$, $\phi=0.85$ is at $(\epsilon_{x},\epsilon_{y})$ location $(0.325,0.177)$, which is far beyond the jamming transition in both extension directions and is even contained within the non-linear strain-stiffening regime. This indicates that any critical packing fraction in the $L$ phase space is not necessarily critical in strain. Discontinuities in the packing fraction data make additional analysis, such as transition packing fraction along the strain sweeps as a function of yarn length, unreliable. These jumps are not the result of the voxel size, shown in S10b. Though there is a maximum voxel side length for reliable volume data (see Fig. S10a), significantly reducing this voxel side length did not result in smoother data. We believe these jumps are the result of small local rearrangements within the stitch. Since the contact potential is very non-linear and highly dependent on the depth of the contact, the compressed volume of the stitch does not meaningfully represent the energy being minimized in the simulation. The total volume of the stitch can vary widely without subsequent jumps in the energy of the stitch. It is possible that a dynamic simulation framework would result in smoother packing fraction data.
Figure S9: Packing fraction as a function of strain for different lengths of yarn per stitch, $L$. Fabrics were extended in both the $x$- (left) and $y$-directions (right). Fabrics that display jammed behavior have a length of yarn per stitch $L\leq 11.5~{}{\rm mm}$. For these simulations, the compression scaling constant is $k=0.1~{}{\rm mN/mm^{2}}$.
Figure S10: (a) Calculated total volume of the stitch within the fabric using the voxel method as a function of voxel side length. Small voxel side lengths ($<0.05$) show stability in the calculated compressed volume of the stitch, whereas larger voxel side lengths show wide variation in volume. For these simulations, the length of yarn per stitch is $L=12.2~{}{\rm mm}$ and the compression scaling constant is $k=0.1~{}{\rm mN/mm^{2}}$. (b) Packing fraction as a function of strain for fabric stretched in the $x$-direction for two different voxel side lengths: $l=0.04$ in blue and $l=0.02$ in black. For these simulations, the length of yarn per stitch is $L=10.5~{}{\rm mm}$ and the compression scaling constant is $k=0.1~{}{\rm mN/mm^{2}}$.
@@
Each contact volume exerts a force on the yarn within the stitch. Individual forces for each contact for a single set of yarn and manufacturing parameters are shown in Fig. S11. The only non-negligible contact forces in the $x$-direction are the same-row contacts and the self-contact. Both of these forces are reduced when the fabric extends in the $x$-direction (Fig. S11a), and thus assist with normal linear elastic behavior when the fabric is pulled in the $x$-direction and inhibit it when the fabric is pulled in the $y$-direction (Fig. S11d). The NN same-column contact and the NNN same-column contact are applying forces in the $y$-direction; the NN same-column contact encourages the fabric to shrink in the $y$-direction while the NNN same-column contact encourages the fabric to grow in the $y$-direction. These forces are stagnant at low strains when the fabric is pulled in the $x$-direction, potentially indicating a lack of contact rearrangement. When the fabric is pulled in the $y$-direction, the force from the NNN same-column contact decreases in magnitude while the NN same-column contact grows. The interplay between these two forces and the forces in the $x$-direction determines jamming in the $y$-direction, as discussed in the Main Text and displayed in Fig. 6h.
@@
When the fabric is stretched in the $x$-direction, the mechanics of jamming show the stitch shrinking out-of-plane (Main Text Fig. 6g). The contact with non-negligible force in this direction is the NN same-column contact, which is applying a force in the direction of expansion, directly contrary to the actual motion of the stitch. This is additional evidence that bending forces dominate jamming in the $x$-direction.
@@
Figure S11: Contact forces as a function of strain in each of the relevant directions for all of the contacts. The sign of the force indicates whether or not it assists with normal linear elastic behavior (positive) or resists linear elastic mechanics (negative) when stretched in that direction. For each type of contact where relevant, both contacts of that type are added (eg: the NN same-column contact is the sum of this contact force with the stitch above and below). Both sets of data are for length of yarn per stitch $L=10.5~{}{\rm mm}$, which is jammed in both directions. (a) Force in the $x$-direction for the fabric stretched in the $x$-direction. The only non-negligible forces are from the same-row (magenta) and self contacts (green), which track together. (b) Force in the $y$-direction for the fabric stretched in the $x$-direction. The NN same-column contact (blue) promotes typical linear elastic extension while the NNN same-column contact (yellow) resists that motion. (c) Force in the $z$-direction (out-of-plane) for the fabric stretched in the $x$-direction. The only contribution in this direction is the NN same-column contact (blue), which is pushing out, contrary to the actual motion of the yarn in this jammed configuration. (d) Force in the $x$-direction for fabric stretched in the $y$-direction. The same-row (magenta) and self (green) contacts are the only contributing contacts, and both are in the direction of expansion. (e) Force in the $y$-direction for fabric stretched in the $y$-direction. The only contributions are the NN same-column (blue) and NNN same-column (yellow) contacts, which resist linear elastic motion and promote it, respectively. The forces in the $z$-direction for fabrics stretched in the $y$-direction are similar to those found when stretching in the $x$-direction, but flipped in sign due to the different out-of-plane linear elastic behavior seen in the $y$-direction (Fig. S12b).
@@
As discussed in the Main Text, jammed fabrics pulled in the $x$-direction have unique mechanics markers in the out-of-plane ($z$) strain. Unjammed fabrics have out-of-plane expansion at low strain and jammed fabrics shrink at low strain. Fig. S12 shows the full range of motion for one unjammed sample and a superjammed sample. In the superjammed sample, the out-of-plane width never recovers to the zero-strain dimension through the entirety of the range of motion. This is due to a lack of a linear elastic regime, which is discussed in more detail in Section S4.
@@
For simulated fabrics pulled in the $x$-direction, the area-per-stitch strain is non-monotonic, first increasing and then beginning to decrease at very large strain as seen in S13. Unlike fabrics stretched in the $y$-direction, there is no marker of jamming in the simulated area strain for fabrics stretched in the $x$-direction. Though we cannot directly compare these results to experimental area strain, seen in S14, because of the “stitch size” affect on yarn radius and potentially the core radius, both the simulations and experiments show periods of increasing area strain followed by decreasing area strain. The simulations also do not consider boundary effects, which would provide large constraints on the cell area, particularly at high strain. These boundary effects may be the cause of the relatively constant area strain at large uniaxial strain, which is not present in simulations.
@@
extension direction $k$ [mN/mm2] $L$ [mm] Step Size [mm] $\epsilon_{final}$ RMS [J]